OpenStudy (anonymous):

By noon on Thursday, Terry had loaded 13 of the company's new laptops with software. If there were 52 new laptops all together, what percentage of the job had he completed

OpenStudy (anonymous):

\[\frac{ 13 }{ 52 }\] computers are done. This simplifies to\[\frac{ 1 }{ 4 }\] and now convert this fraction to a percentage.

OpenStudy (133):

25% because 13 divided by 52=0.25*100=25%

OpenStudy (133):

Hope that helps for future references @Savannaaah